OpenStudy (anonymous):

Which of the following is not a perfect square trinomial? 144 + 12y + y2 100 + 20y + y2 64 − 16y + y2 25 + 10y + y2

OpenStudy (anonymous):

@amistre64

OpenStudy (mathmale):

As an example: 64-16y+y^2: Is 64 a perfect square? If so, what is its square root? Call it a. Is y^2 a perfect square? If so, what is its square root? Call it b. Write out the factors: (a-b)(a+b). Multiply these two binomials together: a^2 -2ab + b^2. If your -2ab is the same as your -16y, then 64-16y+y^2 is a perfect square trinomial. To re-cap: identify your a and your b. Multiply them together. Is their product equal to -16y (except for sign)? If yes, then 64-16y+y^2 is a perfect square. Check the other three trinomials in the same way.

OpenStudy (anonymous):

i don't get it...:(

OpenStudy (amistre64):

expnad out a binomial that is squared. a perfect square is just some value, squared. (y+a)^2 = (y+a)(y+a) expand that out to a polynomial to compare parts

OpenStudy (anonymous):

this is going to take a while for me to learn.

OpenStudy (amistre64):

aa + (a+a)y + y^2 which option does not follow this rule: the factor (a) is doubles to get (a+a)

OpenStudy (amistre64):

start by assessing the sqrts of the constants

OpenStudy (anonymous):

c?

OpenStudy (amistre64):

c means nothing to me, lets work a process. give me a list of the sqrts of the constants

OpenStudy (mathmale):

Compare these two trinomial expressions: aa + (a+a)y + y^2 8*8 + (8+8)y +y^2=64+16y+y^2. (This is one of your answer choices.) What is " a " in this case? Is 64+16y+y^2 a perfect square trinomial?

OpenStudy (anonymous):

yes it is. am i correct?

OpenStudy (amistre64):

there is only one way i would know if you are correct, and that would be to do a simple 3 step process ... step one, take the square roots of the constants, step 2, double the results. step 3 pick the correct answer.

OpenStudy (mathmale):

Since you continue to ask whether your answer is correct or not, I'll take a different approach: Please explain how you decided that " c " is the correct answer.

OpenStudy (anonymous):

I guessed

OpenStudy (anonymous):

I have a lot of stress on me from preparing you PASS testing and its driving me insane

OpenStudy (anonymous):

I have three papers to write and my teachers are breathing down my neck

OpenStudy (anonymous):

and i dont get any of this math

OpenStudy (mathmale):

I spent several minutes trying to explain the problem up front. If you don't understand what I wrote, ask some questions; the answers may help you to understand the problem better.